### v1

@@ 2.5 Diamond investigates enhancer activity in Drosophila embryos
Figure 5: Evaluating Diamond on a real Drosophila enhancer dataset. (a) Diamond is applied with four different ML models to identify important non-additive interactions and is compared against two baseline methods. Each possible interaction is measured by the minimum FDR threshold cutoff at which it is selected, with the top five interaction annotated. The annotated transcription factors are labeled by their UniProt identifiers. The red stars indicate well-characterized interactions in early Drosophila embryos as ground truth. The blue stars indicate interactions supported by literature evidence, referenced by the accompanying PubMed identifiers. The top interactions reported by the MLP model are qualitatively evaluated from three aspects: the contribution of feature values to the (b) marginal and (c) interaction importance measure, and (d) the contribution of the marginal importance measures to the interaction importance measure.
@@
We then evaluated the performance of Diamond on a Drosophila enhancer dataset [3] to investigate the relationship between enhancer activity and DNA occupancy for transcription factor (TF) binding and histone modifications in Drosophila embryos. We used a quantitative study of DNA occupancy for $23$ TFs and $13$ histone modifications with the enhancer status labeled for $7,809$ genomic sequence samples from blastoderm Drosophila embryos (Fig. S.4). The enhancer status for each sequence is binarized as the response, depending on whether the sequence drives patterned expression in blastoderm embryos. To predict the enhancer status, the maximum value of normalized fold-enrichment [27] of ChIP-seq and ChIP-chip assays for each TF or histone modification served as our features. For robustness, we repeated the experiment 20 times using different random seeds.
@@
We started by comparing the identified interactions against a list of well-characterized interactions in the early Drosophila embryos, summarized by [3]. If Diamond is successful in identifying important interactions while controlling the FDR, then the identified interactions should considerably overlap with the ground truth list. Our results show that three of the top five interactions reported by Diamond are present in the ground truth list for MLP and tree-based models, while four out of five interactions are present in the ground truth list for FM (Fig. 5a). In comparison, the baseline methods exhibit significantly different behaviors. One method, which is based on permutation-based interaction-wise p-values and the Benjamini-Hochberg procedure, reported no ground truth interactions among its top five identifications. Another method, which aggregates feature-wise FDR, reported both ground truth along with a large number of non-ground truth interactions.
@@
Next, we investigated the identified interactions that are not included in the ground truth list. For example, we investigated the interaction between the TFs Snail (UniProt ID: P08044) and Twist (UniProt ID: P10627) identified by Diamond+MLP through databases containing assorted experimentally verified interactions [30] and literature [21]. Research indicates that Snail represses Twist targets because their target sequences are very similar, and their binding is mutually exclusive. This literature evidence is further supported by a qualitative evaluation (Fig. 5b-d). Specifically, we examined the Snail–Twist interaction in terms of the contribution of feature values to the enhancer status prediction, observing that high Twist expression suppresses the contribution of Snail to enhancer activation.
@@
Finally, Diamond is able to identify non-additive interactions with various interactive patterns (Fig. 5b-d). Among the top five interactions reported by Diamond with MLP, the Zelda–Twist interaction, the Bicoid–Twist interaction, and the Krueppel–Twist interaction exhibit synergistic effects. The high expression of both TFs in these interactions reinforces the contribution of each individual factor to enhancer activation. In contrast, the Snail–Twist interaction, which was mentioned earlier, exhibits repressive effects, where high expression of one TF suppresses the contribution of the other to enhancer activation. Lastly, the Twist–Zeste interaction exhibits alternative effects, where high expression of either TF enhances the contribution of each to enhancer activation.
@@ 2.6 Diamond understands mortality risk factors in health outcomes
Figure 6: Evaluating Diamond on a real mortality risk dataset. (a) Diamond is applied with the MLP model to identify important non-additive interactions and is compared against other baseline methods. Each possible interaction is measured by the minimum FDR threshold cutoff at which it is selected, with the top five interaction annotated. The blue stars indicate interactions supported by literature evidence, referenced by the accompanying PubMed identifiers. (b) Each feature contributes differently to predicting the mortality status, as measured by the Expected Gradient scores. It is worth mentioning that two marginally important features do not necessarily result in important interactions, as anticipated in Diamond’s design. The top interactions reported by Diamond are qualitatively evaluated from three aspects: the contribution of feature values to the (b) marginal and (c) interaction importance measure, and (d) the contribution of the marginal importance measures to the interaction importance measure.
@@
We lastly evaluated the performance of Diamond on a real mortality risk dataset [10] to investigate the relationship between mortality risk factors and long-term health outcomes within the US population. We used a mortality dataset from the National Health and Nutrition Examination Survey (NHANES I) and NHANES I Epidemiologic Follow-up Study (NHEFS). The dataset incorporates $35$ clinical and laboratory measurements from $14,407$ US participants between 1971 and 1974 (Fig. 6b). Note that some of the features are presented categorically, which makes model’s trained from this data challenging to interpret. We addressed this problem by converting each categorical feature into a list of binary features using one-hot encoding. The dataset also reports the mortality status of participants as of 1992, with $4,785$ individuals known to have passed away before 1992. The task is to construct an MLP-based model to predict mortality status using a survival loss function. For robustness, we repeated the experiment 20 times using different random seeds.
@@
Given the absence of ground truth interactions for this task, we began by evaluating the interactions identified by Diamond through literature support. Our analysis shows that three of the top ten selected interactions are directly supported by existing literature (Fig. 6a). For example, we investigated the interaction between sex and sedimentation rate (SR) and found that, although a high SR is strongly associated with a higher risk of overall mortality, this association can be attenuated by sex [15]. Additionally, we investigated the interaction between creatinine and the blood urea nitrogen (BUN) identified by Diamond. Research indicates that the BUN/creatinine ratio is recognized to have a nonlinear association with all-cause mortality and a linear association with cancer mortality [46]. The literature evidence is further supported by qualitative evaluation (Fig. 5c-e). Specifically, we examined the BUN-creatinine interaction in terms of the contribution of feature values to mortality status prediction, observing that when the creatinine level exceeds a certain level, a high BUN level further increases the mortality risk.
@@ 3 Discussion
In this study, we aim to enable rigorous data-driven scientific discoveries, which are crucial in the present of massive data sets. For this purpose, we propose Diamond, an error-controlled interaction discovery method designed to work with a variety of ML models. The key novelties of Diamond are threefold. First, Diamond achieves FDR control using carefully designed knockoffs, without relying on p-values, which are often difficult to obtain in generic ML models. Second, Diamond includes a calibration procedure to distill non-additive or interactive effects, thereby maintaining FDR control at the target level. This calibration step is critical because naive application of off-the-shelf feature interaction importance measures fail to control the FDR in this setting. Third, Diamond is versatile in discovering important non-additive interactions across a wide range of ML models, different knockoff designs, and various interaction importance measures, all while guaranteeing FDR control. We have applied Diamond to various simulated and real datasets to demonstrate its empirical utility, proving it to be the only valid solution for error-controlled data-driven scientific discovery compared to other alternative methods.
@@
Lastly, this study highlights several promising directions for future research. First, we observe that Diamond tends to be conservative. This suggests a future direction for improvement in enhancing statistical power through further refinement of the FDR estimation process. Previous efforts, like the knockoff-tailored plugin pairwise-coupling layer in MLP models, have shown superior performance in statistical power [33]. However, these previously described methods are specific to certain types of ML models and may not be generally applicable to all generic ML models, indicating the need for improvement to broaden their applicability. Second, we observe that Diamond cannot distinguish between direct interactions and interactions caused by transitive effects. Specifically, even without a direct interaction between two features, we might still expect a strong interaction between them if solid interactions exist individually between each of these features and a third feature. Though transitive interactions are non-additive, for the purpose of scientific discovery, a potential research direction is to identify FDR-controlled direct interactions exclusively, using a causal inference framework [38]. Finally, this study primarily focuses on discovering pairwise interactions. While detecting pairwise interactions holds practical significance in various biological contexts, the discovery of higher-order interactions can naturally provide deeper insights for explaining genetic mechanisms, diseases, and drug effects in healthcare domains. This problem is highly challenging in practice due to the exponentially large search space. A potential research direction for future studies is to effectively reduce the search space to generalize Diamond for practical discovery of higher-order interactions.
In conclusion, Diamond enables error-controlled detection of feature interactions in the context of a variety of ML models. The versatility and flexibility of Diamond make it widely applicable in high-stakes and error-intolerant domains where interpretability and statistical rigor are needed. We believe that this powerful tool will pave the way for the broader deployment of machine learning models in scientific discovery and hypothesis validation.
@@ 4.1 FDR control with the knockoffs
4.2 Measuring non-additive interactive effect
As a key precursor to FDR estimation, Diamond quantifies feature interactions from trained ML models and produces a ranked list of these interactions, with higher-ranked interactions indicating greater importance. For notational simplicity, we use indices for both original features and knockoffs as $\left\{1,2,\cdots,2p\right\}$, with $\left\{1,\cdots,p\right\}$ and $\left\{p+1,\cdots,2p\right\}$ corresponding to the original features and their respective knockoff counterparts. Here, we define $\mathbf{E}^{\text{2D}}=\left[e_{ij}\right]_{i,j=1}^{2p}\in\mathbb{R}^{2p\times 2p}$ as a reported interaction importance measure from existing methods. There are many feature interaction importance measures available for $\mathbf{E}^{\text{2D}}$, each attributing the prediction influence to feature pairs in different ways. However, it is important to note that such measures favor pairs where both features are simultaneously important for a model’s prediction, rather than capturing the true non-additive or interactive effects between the two features [54]. Further supported by simulation studies (Fig. 2c), we observed that many feature interaction importance measures tend to assign higher interaction scores to two marginally important but non-interacting features compared to two random ones, even though neither pair has a real interaction, leading to the failure of FDR control.
The direct reason for the interaction importance measure falling short in controlling FDR is its violation of the knockoff filter’s assumption. Specifically, the knockoff filter requires that the importance scores of knockoff-involving interactions and false interactions have a similar distribution. To resolve this issue, we introduce a calibration procedure to be applied on top of existing interaction importance measures. Specifically, we consider that a reported interaction importance measure from existing methods comprises a mixture of several factors: prediction-dependent marginal effects for individual features, prediction-independent feature biases, independent random noise, and potential non-additive interactive effects between feature pairs. Thus, the reported interaction between features $i$ and $j$ is represented as:
@@ 4.2 Measuring non-additive interactive effect
where $\varepsilon_{ij}\in\mathbb{R}$ is random noise independent of both features and predictions, $e_{ij}\in\mathbb{R}$ and $e_{i},e_{j}\in\mathbb{R}$ are reported pairwise and univariate feature importance measures that are dependent on the model’s predictions, respectively. The functions $g_{i},g_{j}:\mathbb{R}\mapsto\mathbb{R}$ adapt univariate feature importance to be compatible with feature interaction importance. The function $b:\mathbb{R}^{2p}\mapsto\mathbb{R}$ models the feature-specific biases that are independent of the model’s predictions, where $I_{ij}\in\left\{0,1\right\}^{2p}$ indicates the presence of feature $i$ and $j$. Our goal is to identify $s_{ij}$, the potential non-additive interactive effects between features $i$ and $j$.
We formulate the identification of interactive effects $s_{ij}$ as the residuals of a regression task:
@@ 4.3 FDR control for interactions
After calculating the non-additive interactive effects using Eq. 4, we denote the resultant set of interactive effects as $\Gamma=\left\{s_{ij}|i<j,i\neq j-p\right\}$. We arrange $\Gamma$ in decreasing order and select interactions for which the interactive effect, $\Gamma_{j}$, exceeds some threshold, $T$. This selection ensures that the chosen interactions adhere to a desired FDR level $q\in(0,1)$.
@@
where $\mathcal{D}$ and $\mathcal{DD}$ respectively denote the sets of interactions that include at least one knockoff feature and both knockoff features, while $\mathcal{T}$ refers to the set of unique nonzero values present in $\Gamma$.
4.4 ML models and interpretation

### v2

@@ 2.5 Diamond investigates enhancer activity in Drosophila embryos
Figure 5: Evaluating Diamond on a real Drosophila enhancer dataset. (a) Each feature contributes differently to predicting enhancer status, as measured by the Expected Gradient scores in the MLP model. (b) Diamond is compared against three baseline methods. The annotated transcription factors are labeled by their UniProt identifiers. The red stars indicate well-characterized physical interactions in early Drosophila embryos as ground truth. The blue stars indicate interactions supported by literature evidence, referenced by the accompanying PubMed identifiers. (c) Diamond is used with various ML models to identify important non-additive interactions. Each possible interaction is measured by the minimum FDR threshold cutoff at which it is selected, with the top five interaction annotated. The top interactions reported by the MLP model are qualitatively evaluated from three aspects: the contribution of feature values to the (d) marginal and (e) interaction importance measure, and (f) the contribution of the marginal importance measures to the interaction importance measure.
@@
We then evaluated the performance of Diamond on a Drosophila enhancer dataset [4] to investigate the relationship between enhancer activity and DNA occupancy for transcription factor (TF) binding and histone modifications in Drosophila embryos. We used a quantitative study of DNA occupancy for $23$ TFs and $13$ histone modifications with the enhancer status labeled for $7,809$ genomic sequence samples from blastoderm Drosophila embryos (Fig. 4a). The enhancer status for each sequence is binarized as the response, depending on whether the sequence drives patterned expression in blastoderm embryos. To predict the enhancer status, the maximum value of normalized fold-enrichment [31] of ChIP-seq and ChIP-chip assays for each TF or histone modification served as our features. We considered different types of ML models: MLP, FT-Transformer, KAN, LightGBM, XGBoost, and FM. We assessed the trained ML model using 5-fold cross-validation and selected the one with the best performance for model interpretation and interaction discovery. For robustness, we repeated the experiment 20 times using different random seeds. Each repetition involved knockoff generation, ML model training, and interaction-wise FDR estimation.
@@
We started by comparing the identified interactions against a list of well-characterized physical interactions in the early Drosophila embryos, collected by [4]. These interactions have been identified over decades and have been confirmed to play a critical role in regulating spatial and temporal patterning [47]. If Diamond is successful in identifying important interactions while controlling the FDR, then the identified interactions should considerably overlap with these previously reported physical interactions. Meanwhile, it is important to acknowledge that this list of interactions, proposed years ago, represents only a partial set of the true interactions. Therefore, a detected interaction not appearing in the list does not necessarily indicate it is incorrect. Our results show that three of the top five interactions reported by Diamond are included in the ground truth list for MLP and tree-based models, four out of five for FM, and all five for the FT-Transformer (Fig. 5c). The only exception is KAN, which reported all but one interactions at the same FDR levels. In comparison, the baseline methods exhibit significantly different behaviors (Fig. 5b). Two baseline methods, which are based on permutation-based interaction-wise p-values combined with the Benjamini-Hochberg and the Benjamini-Yekutieli procedures respectively, reported no ground truth interactions among its top five identifications. Another method, which aggregates feature-wise FDR, reported ground truth interactions along with a large number of non-ground truth interactions.
@@
Next, we investigated the identified interactions that are not included in the ground truth list. For example, we investigated the interaction between the TFs Snail (UniProt ID: P08044) and Twist (UniProt ID: P10627) identified by Diamond+MLP through databases containing assorted experimentally verified interactions [34] and literature [24]. Research indicates that Snail represses Twist targets because their target sequences are very similar, and their binding is mutually exclusive. This literature evidence is further supported by a qualitative evaluation (Fig. 5d-f, Fig. A.4, Fig. A.5, and Fig. A.6). Specifically, we examined the Snail–Twist interaction in terms of the contribution of feature values to the enhancer status prediction, observing that high Twist expression suppresses the contribution of Snail to enhancer activation.
@@
Finally, Diamond is able to identify non-additive interactions with various interactive patterns (Fig. 5d-f). Among the top five interactions reported by Diamond with MLP, the Zelda–Twist interaction, the Bicoid–Twist interaction, and the Krueppel–Twist interaction exhibit synergistic effects. The high expression of both TFs in these interactions reinforces the contribution of each individual factor to enhancer activation. In contrast, the Snail–Twist interaction, which was mentioned earlier, exhibits repressive effects, where high expression of one TF suppresses the contribution of the other to enhancer activation. Lastly, the Twist–Zeste interaction exhibits alternative effects, where high expression of either TF enhances the contribution of each to enhancer activation.
@@ 2.6 Diamond understands mortality risk factors in health outcomes
Figure 6: Evaluating Diamond on a real mortality risk dataset. (a) Each feature contributes differently to predicting the mortality status, as measured by the Expected Gradient scores. It is worth mentioning that two marginally important features do not necessarily result in important interactions, as anticipated in Diamond’s design. (b) Diamond is used with the MLP model to identify important non-additive interactions and is compared against other baseline methods. Each possible interaction is measured by the minimum FDR threshold cutoff at which it is selected, with the top five interaction annotated. The blue stars indicate interactions supported by literature evidence, referenced by the accompanying PubMed identifiers. The survival loss function for predicting mortality restricts the modeling to DNNs, and we use MLP specifically due to memory limitations with other DNN models. The top interactions reported by Diamond are qualitatively evaluated from three aspects: the contribution of feature values to the (c) marginal and (d) interaction importance measure, and (e) the contribution of the marginal importance measures to the interaction importance measure.
@@
We lastly evaluated the performance of Diamond on a real mortality risk dataset [12] to investigate the relationship between mortality risk factors and long-term health outcomes within the US population. We used a mortality dataset from the National Health and Nutrition Examination Survey (NHANES I) and NHANES I Epidemiologic Follow-up Study (NHEFS). The dataset incorporates $35$ clinical and laboratory measurements from $14,407$ US participants between 1971 and 1974 (Fig. 6a). Note that some of the features are presented categorically, which makes model’s trained from this data challenging to interpret. We addressed this problem by converting each categorical feature into a list of binary features using one-hot encoding. The dataset also reports the mortality status of participants as of 1992, with $4,785$ individuals known to have passed away before 1992. The task is to construct an MLP-based model to predict mortality status using a survival loss function. We assessed the trained model using 5-fold cross-validation and selected the one with the best performance for model interpretation and interaction discovery. For robustness, we repeated the experiment 20 times using different random seeds. Each repetition involved knockoff generation, ML model training, and interaction-wise FDR estimation.
@@
Given the absence of ground truth interactions for this task, we began by evaluating the interactions identified by Diamond through literature support. Our analysis shows that three of the top ten selected interactions are directly supported by existing literature (Fig. 6b). For example, we investigated the interaction between sex and sedimentation rate (SR) and found that, although a high SR is strongly associated with a higher risk of overall mortality, this association can be attenuated by sex [17]. Additionally, we investigated the interaction between creatinine and the blood urea nitrogen (BUN) identified by Diamond. Research indicates that the BUN/creatinine ratio is recognized to have a nonlinear association with all-cause mortality and a linear association with cancer mortality [52]. The literature evidence is further supported by qualitative evaluation (Fig. 6c-e). Specifically, we examined the BUN-creatinine interaction in terms of the contribution of feature values to mortality status prediction, observing that when the creatinine level exceeds a certain level, a high BUN level further increases the mortality risk. In contrast, the baseline methods display significantly different behavior, reporting an overwhelming number of interactions at the significant FDR level, which appears to overestimate false positives.
@@ 3 Discussion
In this study, we aim to enable rigorous data-driven scientific discoveries, which are crucial in the present of massive data sets. For this purpose, we propose Diamond, an error-controlled interaction discovery method designed to work with a variety of ML models. The key novelties of Diamond are threefold. First, Diamond achieves FDR control using carefully designed knockoffs, without relying on p-values, which are often difficult to obtain in generic ML models. Second, Diamond distills non-additive interaction effects, thereby maintaining FDR control at the target level. This non-additivity distillation step is critical because naive application of off-the-shelf feature interaction importance measures fail to control the FDR in this setting. Third, Diamond is versatile in discovering important non-additive interactions across a wide range of ML models, different knockoff designs, and various interaction importance measures, all while guaranteeing FDR control. We have applied Diamond to various simulated and real datasets to demonstrate its empirical utility, proving it to be the only valid solution for error-controlled data-driven scientific discovery compared to other alternative methods.
@@
Lastly, this study highlights several promising directions for future research. First, we observe that Diamond tends to be conservative. This suggests a future direction for improvement in enhancing statistical power through further refinement of the FDR estimation process. Previous efforts, like the knockoff-tailored plugin pairwise-coupling layer in MLP models, have shown superior performance in statistical power [38]. However, these previously described methods are specific to certain types of ML models and may not be generally applicable to all generic ML models, indicating the need for improvement to broaden their applicability.
Secondly, we observe that Diamond cannot distinguish between direct interactions and interactions caused by transitive effects. In reality, the former is scientifically more interesting than the latter, though both fit the definition of non-additive interactions. Specifically, even without a direct interaction between two features, we might still expect a strong interaction between them if solid interactions exist individually between each of these features and a third feature. Though transitive interactions are non-additive, for the purpose of scientific discovery, a potential research direction is to identify FDR-controlled direct interactions exclusively, using a causal inference framework [43]. Thirdly, we observe that different ML models exhibit different levels of performance in identifying important interactions. Briefly, models with stronger modeling capabilities, such as transformer models, are better at identifying important interactions compared to models with weaker capabilities, such as FMs and CNNs. As independent studies have highlighted, the sensitivity of interaction detection depends on the importance measures of the interactions, which are critically influenced by the quality of the model [1]. A potential research direction is to automatically determine the best ML models for interaction detection, rather than focusing solely on modeling accuracy.
Finally, this study primarily focuses on discovering pairwise interactions. While detecting pairwise interactions holds practical significance in various biological contexts, the discovery of higher-order interactions can naturally provide deeper insights for explaining genetic mechanisms, diseases, and drug effects in healthcare domains. This problem is highly challenging in practice due to the exponentially large search space. A potential research direction for future studies is to effectively reduce the search space to generalize Diamond for practical discovery of higher-order interactions.
In conclusion, Diamond enables error-controlled detection of non-additive feature interactions in the context of a variety of ML models. The versatility and flexibility of Diamond make it widely applicable in high-stakes and error-intolerant domains where interpretability and statistical rigor are needed. The robustness against misspecified ML models and low-quality knockoffs ensures the reliability of Diamond, even when erroneously applied in extreme cases. We believe that this powerful tool will pave the way for the broader deployment of machine learning models in scientific discovery and hypothesis validation.
@@ 4.1 FDR control with the knockoffs
4.2 Measuring non-additive interaction effect
As a key precursor to FDR estimation, Diamond quantifies feature interactions from trained ML models and produces a ranked list of these interactions, with higher-ranked interactions indicating greater importance. For notational simplicity, we use indices for both original features and knockoffs as $\left\{1,2,\cdots,2p\right\}$, with $\left\{1,\cdots,p\right\}$ and $\left\{p+1,\cdots,2p\right\}$ corresponding to the original features and their respective knockoff counterparts. Here, we define $\mathbf{E}^{\text{2D}}=\left[e_{ij}\right]_{i,j=1}^{2p}\in\mathbb{R}^{2p\times 2p}$ as a reported interaction importance measure from existing methods. There are many feature interaction importance measures available for $\mathbf{E}^{\text{2D}}$, each attributing the prediction influence to feature pairs in different ways. However, it is important to note that such measures favor pairs where both features are simultaneously important for a model’s prediction, rather than capturing the true non-additive interaction effects between the two features [61]. Further supported by simulation studies (Fig. 2c), we observed that many feature interaction importance measures tend to assign higher interaction scores to two marginally important but non-interacting features compared to two random ones, even though neither pair has a real interaction, leading to the failure of FDR control.
The direct reason for the interaction importance measure falling short in controlling FDR is its violation of the knockoff filter’s assumption. Specifically, the knockoff filter requires that the importance scores of knockoff-involving interactions and false interactions have a similar distribution. To resolve this issue, we introduce a non-additivity distillation procedure to be applied on top of existing interaction importance measures. Specifically, we consider that a reported interaction importance measure from existing methods comprises a mixture of several factors: label-dependent marginal effects for individual features, label-independent feature biases, independent random noise, and potential non-additive interaction effects between feature pairs. Thus, the reported interaction between features $i$ and $j$ is represented as:
@@ 4.2 Measuring non-additive interaction effect
where $\varepsilon_{ij}\in\mathbb{R}$ is random noise independent of both features and predictions, $e_{ij}\in\mathbb{R}$ and $e_{i},e_{j}\in\mathbb{R}$ are reported pairwise and univariate feature importance measures that are dependent on the model’s predictions, respectively. The functions $g_{i},g_{j}:\mathbb{R}\mapsto\mathbb{R}$ adapt univariate feature importance to be compatible with feature interaction importance. The function $b:\mathbb{R}^{2p}\mapsto\mathbb{R}$ models the feature-specific biases that are independent of the model’s predictions, where $I_{ij}\in\left\{0,1\right\}^{2p}$ indicates the presence of feature $i$ and $j$. Our goal is to identify $s_{ij}$, the potential non-additive interaction effects between features $i$ and $j$.
We formulate the identification of interaction effects $s_{ij}$ as the residuals of a regression task:
@@ 4.3 FDR control for interactions
After calculating the non-additive interaction effects using Eq. 4, we denote the resultant set of interaction effects as $\Gamma=\left\{s_{ij}|i<j,i\neq j-p\right\}$. We arrange $\Gamma$ in decreasing order and select interactions for which the interaction effect, $\Gamma_{j}$, exceeds some threshold, $T$. This selection ensures that the chosen interactions adhere to a desired FDR level $q\in(0,1)$.
@@
where $\mathcal{K}$ and $\mathcal{KK}$ respectively denote the sets of interactions that include at least one knockoff feature and both knockoff features, while $\mathcal{T}$ refers to the set of unique nonzero values present in $\Gamma$. Refer to Sec. A.4 for more detailed description of FDR estimation.
4.4 ML models